OpenStudy (anonymous):

what is the y-coordinate for the point of intersection of the graphs of 2x + 3y = -6 and 3x + 4y =5?

OpenStudy (anonymous):

Well, if you want the equations to intersect you simply set them equal. So, solve them both for y. Set them equal, solve for x. Then, once you find x, plug back in to solve for y. You only need to plug it into one equation because you know its an intersection point so it has to have the same y coordinate!! Let me know if you need more help

OpenStudy (anonymous):

i still need help so once you solve them both for y, can you show me where you go from there?=)

OpenStudy (anonymous):

2x+3y=-6,.(1) 3x+4y=5....(2) (1)*3-(2)*2 gives 9y-8y=-18-10 y=-28

OpenStudy (anonymous):

Sure. Solving the first one for y you get y=-2-(2/3)x and the second is y=(5/4)-(3/4)x. Set them equal you get: -2-(2/3)x=(5/4)-(3/4)x Add the 2 and add the (3/4)x to get: (13/4)=(1/12)x then multiply by 12 to get 156/4=x or 114=x. Then plug that in. Or do elimination as dip did.

OpenStudy (anonymous):

wait how did you get y=-2 -2/3x and y=5/4 -3/4x? can you explain that?

OpenStudy (anonymous):

can you explain what you did in like the first 4 steps dipankarstudy??? sorry im not very could at math so use simple words..... thx=)

OpenStudy (anonymous):

All I did was solve for y. So I subtracted the 2x from the first one then divided by 3 which is in front of the y. I did a similar process for the other one. Dip multiplied the first equation by 3 and the second one by 2 and then added them. This is called elimination. It allows you to take a multiple of each equation and add them to cancel out one of the variables. In this case, it cancels out the x and you can solve directly for y without the intermediate step that I had. My method is called substitution. Both are equally valid.

OpenStudy (anonymous):

thx i got it...